OpenStudy (anonymous):

Show that function F(x) = x sin x is an antiderivative of f(x) = sin x +x cos x. Determine the antiderivative of f(x) which passes point (0, 3).

OpenStudy (amistre64):

show? as in derive by the rules or by first principles?

OpenStudy (anonymous):

Also, how do you correctly say "pass a point" mathematically in english?

OpenStudy (anonymous):

By first principles, please.

OpenStudy (amistre64):

passes thru the point im assuming is what it means

OpenStudy (anonymous):

Right, thanks for that!

OpenStudy (amistre64):

first principles are a pain with trig becasue there is alot of hand waiving involved; like a magician or a hypnotist .... its true because i say so type thing

OpenStudy (amistre64):

there is no good way to prove sin(x)/x goes to 1 as x goes to 0 :)

OpenStudy (anonymous):

Right, well then I can just derive F(x) and prove it that way.

OpenStudy (amistre64):

\[\lim_{h ->0}\frac{sin(x+h)-sin(x)}{h}\] to derive by first principles is still deriving :) the "rules" provide established shortcuts

OpenStudy (anonymous):

Heh, I understand. :)

OpenStudy (anonymous):

Would it be ok to integrate f(x) instead?

OpenStudy (anonymous):

This is actually really easy, thanks for the help though. :)

OpenStudy (amistre64):

but for practice: \[\frac{sin(x)cos(h)+sin(h)cos(x)-sin(x)}{h}\] \[\frac{sin(x)(cos(h)-1)+sin(h)cos(x)}{h}\] \[\frac{sin(x)(cos(h)-1)}{h}+\frac{sin(h)cos(x)}{h}\] \[\lim_{h->0}\left( sin(x)\frac{(cos(h)-1)}{h}+cos(x)\frac{sin(h)}{h}\right)\] and for the hand waiving part: [cos(h)-1]/h -> 0 ; and sin(h)/h -> 1 leaveing us with ... [sin(x)]' = cos(x) ... tada lol

OpenStudy (anonymous):

haha, thanks! how do you break the LaTeX lines like you did there?

OpenStudy (amistre64):

integrate [sin(x)+xcos(x)] and see if it produces [x sin(x)] should be prove enough

OpenStudy (amistre64):

using the delimiters [ ] make seperate lines.

OpenStudy (amistre64):

\[stuff in here\*] ^ remove * to format

OpenStudy (anonymous):

\[\pi = 3\] \[0 = 0\]

OpenStudy (sriram):

can u integrate x*cosx

OpenStudy (amistre64):

for inline; change the [ ] to ( )

OpenStudy (amistre64):

spose im typing along and I want it in the sentence: \(\pi \ r^2\). then I use the ( ) delimiters

OpenStudy (anonymous):

Like this \(\pi = 3\)? Awesome!

OpenStudy (amistre64):

lol; and if you want to get real fancy: \(\color{purple}{\text{just add some color}}\) :)

OpenStudy (anonymous):

Thanks so much! I'm very new to LaTeX. :)

OpenStudy (anonymous):

Sriram, you can give me an example of how, if you want a medal. :)

OpenStudy (amistre64):

\begin{array}l\color{red}{\normalsize\text{T}}\color{orange}{\normalsize\text{a}}\color{#9c9a2e}{\normalsize\text{s}}\color{green}{\normalsize\text{t}}\color{blue}{\normalsize\text{e}}\color{purple}{\normalsize\text{ }}\color{purple}{\normalsize\text{t}}\color{red}{\normalsize\text{h}}\color{orange}{\normalsize\text{e}}\color{#9c9a2e}{\normalsize\text{ }}\color{green}{\normalsize\text{r}}\color{blue}{\normalsize\text{a}}\color{purple}{\normalsize\text{i}}\color{purple}{\normalsize\text{n}}\color{red}{\normalsize\text{b}}\color{orange}{\normalsize\text{o}}\color{#9c9a2e}{\normalsize\text{w}}\color{green}{\normalsize\text{!}}\color{blue}{\normalsize\text{!}}\color{purple}{\normalsize\text{ }}\color{purple}{\normalsize\text{l}}\color{red}{\normalsize\text{o}}\color{orange}{\normalsize\text{l}}\color{#9c9a2e}{\normalsize\text{}}\end{array}

OpenStudy (anonymous):

\[\color{red}{test}\]

OpenStudy (anonymous):

\color{yourcolor}{yourtext} So for instance: \color{red}{yourtext}

OpenStudy (anonymous):

Actually you need to write the text outside of equation mode \color{yourcolor}{\text{yourtext}} \[\color{blue}{\text{This is is the result}}\]

OpenStudy (anonymous):

Yea, I noticed I can study the LaTeX source by right clicking on the text and choosing 'show source'. :)